# Holon metal, charge-density-wave and chiral superconductor from doping fractional Chern insulator and SU(3)1 chiral spin liquid

### v1

@@ Appendix B Topological property of superconductor from CDW metal
In this case we have a $p\pm ip$ SC coexisting with $U(1)_{2}$ topological order. The chiral central charge is $c=1\pm\frac{1}{2}$. Starting from the theory for the holon metal in Eq. 24, a non-zero s wave $\Delta_{12}$ can be described by a simple term $-\frac{1}{2\pi}ad(\alpha+\beta)$, which then locks $\beta=-\alpha$. In the end the CDW metal is simply a Fermi liquid of $\psi_{3}$ plus a $U(1)_{2}$ topological order described by $\frac{2}{4\pi}\alpha d\alpha$. Note that this topological order does not couple to $A$ and only contributes a chiral central charge of $c=1$. If we further introduce $p\pm ip$ pairing for $\psi_{3}$, we simply have a $p\pm ip$ superconductor coexisting with a hidden topological order with $2$ anyons.
@@ B.2 $\Delta_{12}$ in $p+ip$ channel
We assume a $p+ip$ pairing for $\Delta_{12}$. And also a $p\pm ip$ pairing for $\Delta_{33}$. $\Delta_{12}$ gives a term $-\frac{1}{4\pi}(a_{2}-a_{3})d(a_{2}-a_{3})-\frac{1}{2\pi}a_{2}d\alpha-\frac{1}{2\pi}a_{3}d\beta$, where $a_{2},a_{3}$ are introduced to represent the dual field of the pairing $\Delta_{12}$ and also the topological response. $\Delta_{33}$ gives a term $\frac{2}{2\pi}a_{1}d(-A-\alpha-\beta)$. One should keep in mind that the unit charge of $a_{1}$ binds a Majorana zero mode.
@@
Now the charge vector is $\tilde{q}=(-2,0,0,1,1)^{T}$. One can check that $\det W=1$ so the charge quantization for $\tilde{a}$ is not altered. We label the first gauge field for $\tilde{K}$ as $\tilde{a}_{1}$, then there is a term $-\frac{2}{2\pi}Ad\tilde{a}_{1}$, indicating a charge $2e$ superconductor, which coexists with a decoupled sector described by $K^{\prime}$. One can check that the unit charge of $a_{1}$ also carries unit charge of $\tilde{a}_{1}$ and vice versa. Hence the unit charge of $\tilde{a}_{1}$, the elementary vortex of the superconductor, binds a Majorana zero mode. Meanwhile $|\det K^{\prime}|=1$, so there is no deconfined anyon. The final phase is basically just a conventional $p\pm ip$ superconductor. But there is a chiral central charge of $2$ from $K^{\prime}$, so the final central charge is $c=3\pm\frac{1}{2}$ for $p\pm ip$ pairing of $\psi_{3}$.
@@ Appendix C Superconductor directly from the holon metal
One also find that in the new basis, the charge under $\gamma$ is $\tilde{l}_{\gamma}=l_{1}+l_{2}+l_{3}=V_{\mathrm{Ising}}$. The unit charge of $\gamma$ represents the vortex of this charge $2e$ superconductor. Thefore for each vorex hosts a Majorana mode mod $2$ and single Majorana mode only exists in the superconductor vortex. We have $\det K^{\prime}=16$ and the associated chiral central charge is 00. The three $p+ip$ pairings provide a net chiral central charge of $c=3$. Therefore the final phase is topologically equivalent to three copies of $p+ip$ superconductor stacked with an Abelian topological order with $16$ anyons. We note that the $\frac{1}{2\pi}Ad(\beta_{3}+\beta_{4})$ term indicates that the associated Abelian order has anyons with fractional charge. The minimal charge is carried by an anti-semion with charge $e/2$.
@@ Appendix D Superconductor when doping $Q=\frac{2}{3}e$ anyons
Note that $\Phi_{I}$ carries charge $2$ under $a_{\mu}$, so it creates a $2/3$ anyon. At physical density $n=-\frac{2}{3}-x$, we have $n_{\Phi_{I}}=\frac{x}{2}$. From the variation of $a_{0}$ we find that $\frac{\langle da\rangle}{2\pi}=-\frac{2}{3}\sum_{I}n_{\Phi_{I}}=-x$. Each $\Phi_{I}$ is thus at effective magnetic filling of $\nu_{m}=-\frac{1}{4}$. We can simply put each $\Phi_{I}$ in a $-1/4$ Laughlin state.
@@
The charge of the new gauge field is related to the old ones by $q_{\tilde{a}}=-2q_{a}+\sum_{I}q_{\beta_{I}}$ and $q_{\tilde{\beta}_{I}}=q_{\beta_{I}}$. Define $\tilde{l}=(q_{\tilde{a}},q_{\tilde{\beta}_{1}},q_{\tilde{\beta}_{2}},q_{\tilde{\beta}_{3}})$ and $l=(q_{a},q_{\beta_{1}},q_{\beta_{2}},q_{\beta_{3}})$. One can see that $\tilde{l}=(1,1,0,0)$ is allowed by $l=(0,1,0,0)$. Therefore the minimal charge of $\tilde{a}$ is $1$ and this creates the elementary vortex of the charge $2e$ superconductor. It has a statistics of $\theta=-\frac{\pi}{4}$. Because it carries a vortex charge, it costs infinite energy and should not be identified as a deconfined anyon. There is an elementary anyon with $\tilde{l}=(0,1,1,0)$, corresponding to $l=(1,1,1,0)$. It has statistics of $\theta=-\frac{\pi}{2}$ and is an anti-semion. Including the trivial anyon, there are in total two deconfined anyons. The chiral central charge of this superconductor is $c=-2$ if we also include the background integer quantum Hall part from the $\nu=-1$ state.

### v2

@@ Appendix B Topological property of superconductor from CDW metal
In this case we have a $p\pm ip$ SC coexisting with $U(1)_{2}$ topological order. The chiral central charge is $c=1\pm\frac{1}{2}$. Starting from the theory for the holon metal in Eq. 24, a non-zero s wave $\Delta_{12}$ can be described by a simple term $\frac{1}{2\pi}ad(\alpha+\beta)$, which then locks $\beta=-\alpha$. In the end the CDW metal is simply a Fermi liquid of $\psi_{3}$ plus a $U(1)_{2}$ topological order described by $\frac{2}{4\pi}\alpha d\alpha$. Note that this topological order does not couple to $A$ and only contributes a chiral central charge of $c=1$. If we further introduce $p\pm ip$ pairing for $\psi_{3}$, we simply have a $p\pm ip$ superconductor coexisting with a hidden topological order with $2$ anyons.
@@ B.2 $\Delta_{12}$ in $p+ip$ channel
We assume a $p+ip$ pairing for $\Delta_{12}$. And also a $p\pm ip$ pairing for $\Delta_{33}$. $\Delta_{12}$ gives a term $\frac{1}{4\pi}(a_{2}-a_{3})d(a_{2}-a_{3})-\frac{1}{2\pi}a_{2}d\alpha-\frac{1}{2\pi}a_{3}d\beta$, where $a_{2},a_{3}$ are introduced to represent the dual field of the pairing $\Delta_{12}$ and also the topological response. $\Delta_{33}$ gives a term $\frac{2}{2\pi}a_{1}d(-A-\alpha-\beta)$. One should keep in mind that the unit charge of $a_{1}$ binds a Majorana zero mode.
@@
Now the charge vector is $\tilde{q}=(-2,0,0,1,1)^{T}$. One can check that $\det W=1$ so the charge quantization for $\tilde{a}$ is not altered. We label the first gauge field for $\tilde{K}$ as $\tilde{a}_{1}$, then there is a term $\frac{2}{2\pi}Ad\tilde{a}_{1}$, indicating a charge $2e$ superconductor, which coexists with a decoupled sector described by $K^{\prime}$. One can check that the unit charge of $a_{1}$ also carries unit charge of $\tilde{a}_{1}$ and vice versa. Hence the unit charge of $\tilde{a}_{1}$, the elementary vortex of the superconductor, binds a Majorana zero mode. Meanwhile $|\det K^{\prime}|=1$, so there is no deconfined anyon. The final phase is basically just a conventional $p\pm ip$ superconductor. But there is a chiral central charge of $2$ from $K^{\prime}$, so the final central charge is $c=3\pm\frac{1}{2}$ for $p\pm ip$ pairing of $\psi_{3}$.
@@ Appendix C Superconductor directly from the holon metal
One also find that in the new basis, the charge under $\gamma$ is $\tilde{l}_{\gamma}=l_{1}+l_{2}+l_{3}=V_{\mathrm{Ising}}$. The unit charge of $\gamma$ represents the vortex of this charge $2e$ superconductor. Thefore for each vorex hosts a Majorana mode mod $2$ and single Majorana mode only exists in the superconductor vortex. We have $\det K^{\prime}=16$ and the associated chiral central charge is 0. The three $p+ip$ pairings provide a net chiral central charge of $c=3$. Therefore the final phase is topologically equivalent to three copies of $p+ip$ superconductor stacked with an Abelian topological order with $16$ anyons. We note that the $\frac{1}{2\pi}Ad(\beta_{3}+\beta_{4})$ term indicates that the associated Abelian order has anyons with fractional charge. The minimal charge is carried by an anti-semion with charge $e/2$.
@@ Appendix D Superconductor when doping $Q=\frac{2}{3}e$ anyons
Note that $\Phi_{I}$ carries charge $2$ under $a_{\mu}$, so it creates a $2/3$ anyon. At physical density $n=-\frac{2}{3}-x$, we have $n_{\Phi_{I}}=\frac{x}{2}$. From the variation of $a_{0}$ we find that $\frac{\langle da\rangle}{2\pi}=-\frac{2}{3}\sum_{I}n_{\Phi_{I}}=-x$. Each $\Phi_{I}$ is thus at effective magnetic filling of $\nu_{m}=-\frac{1}{4}$. We can simply put each $\Phi_{I}$ in a 1/4-1/4 Laughlin state.
@@
The charge of the new gauge field is related to the old ones by $q_{\tilde{a}}=-2q_{a}+\sum_{I}q_{\beta_{I}}$ and $q_{\tilde{\beta}_{I}}=q_{\beta_{I}}$. Define $\tilde{l}=(q_{\tilde{a}},q_{\tilde{\beta}_{1}},q_{\tilde{\beta}_{2}},q_{\tilde{\beta}_{3}})$ and $l=(q_{a},q_{\beta_{1}},q_{\beta_{2}},q_{\beta_{3}})$. One can see that $\tilde{l}=(1,1,0,0)$ is allowed by $l=(0,1,0,0)$. Therefore the minimal charge of $\tilde{a}$ is $1$ and this creates the elementary vortex of the charge $2e$ superconductor. It has a statistics of $\theta=-\frac{\pi}{4}$. Because it carries a vortex charge, it costs infinite energy and should not be identified as a deconfined anyon. There is an elementary anyon with $\tilde{l}=(0,1,1,0)$, corresponding to $l=(1,1,1,0)$. It has statistics of $\theta=-\frac{\pi}{2}$ and is an anti-semion. Including the trivial anyon, there are in total two deconfined anyons. The chiral central charge of this superconductor is c=2c=-2 if we also include the background integer quantum Hall part from the $\nu=-1$ state.